Circular No. 7655

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001cy IN UGC 11927 M. Ganeshalingam, M. Modjaz, and W. D. Li, University of Califor-

nia at Berkeley, report the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on June 30.5 (mag ∼ 16.3) and July 1.4 UT (mag ∼ 16.2). The new object is located at α = 22^{h}09^{m}27^{s}.76, δ = +40^{o}59′16′′.5 (equinox 2000.0), which is 3′′.6 east and 6′′.1 south of the nucleus of UGC 11927. A KAIT image taken on June 19.5 showed nothing at this position (limiting mag ∼ 19.5).

COMETS C/2001 M7, C/2001 M8, C/2001 M9 (SOHO) Further to IAUC 7650, D. Biesecker and D. Hammer report measure-

ments of three additional Kreutz sungrazing comets found by K. Černis, X. Leprette, and M. Oates on SOHO website images. C/2001 M7 and C/2001 M8 were visible in both the C2 and C3 instruments, while C/2001 M9 was visible only in C2 images. The reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below.

Comet  2001 UT  α_{2000}  δ_{2000}  MPEC
C/2001 M7 June 25.613  6^{h}09.^{m}6  +21^{o}19′  2001-M44
C/2001 M8  26.576  6 15.1  +20 21  2001-M46
C/2001 M9  26.663  6 16.8  +21 44  2001-M55

COMET C/2001 M10 (NEAT) M. D. Hicks, Jet Propulsion Laboratory, reports that this comet shows

a diffuse coma of diameter ∼ 5′′ and a faint tail ∼ 7′′ long in p.a. 240^{o} in a 10-min R-band CCD exposure obtained on June 30.3 UT with the 0.61m reflector at Table Mountain (observers D. Esqueda, Hicks, and T. Ha). Hicks’ name also should be added to the list of NEAT team members on IAUC 7654.

V1178 SCORPII Additional broadband V magnitudes by W. Liller (cf. IAUC 7650):

June 24.006 UT, 11.35; 24.101, 11.36; 24.145, 11.35; 24.215, 11.36; 25.979, 12.27; 26.010, 12.33; 26.109, 12.23; 26.192, 12.22; 26.978, 11.70; 27.007, 11.70; 27.106, 11.85; 27.184, 11.74.

2001 July 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green